Calculate the mass of urea required in making 2.5kg of 0.25molal aqueous solution

Grade:12

1 Answers

Arun
4 years ago
 
Molality(m) = 0.25 mol kg-1
mass of the solution = 2.5 kg = 2500 g
To find,
mass of urea required = ?
 
[Molality = given mass of solute/molar mass × 1000/mass of the solvent (in g)]
 
molar mass=(14+2+12+16+14+2) =60 g mol-1 
Let x be the mass of urea required.
mass of the solvent = (2500-x)
substituting the values,
0.25=x/60 ×1000/(2500-x)
(2500-x)=x/60 ×1000/0.25
(2500-x)=200x/3
2500=203x/3
x=7500/203
x=36.946 g